OpenStudy (anonymous):

A bank contains 30 coins, consisting of nickels, dimes, and quarters. yjere are twice as many nickels as quarters and the remaining coins are dimes. if the total value of the coins is $3.35, what is the number of each type of coin in the bank?

OpenStudy (radar):

Let x = the number of quarters then 2x = the numbers of nickels (given) Then 30 - (x + 2x) are the number of dimes. This simplifies to: 30 - 3x Note this is the number of coins (not the value) Now for the money!

OpenStudy (radar):

25x is the value of the quarters 10x is the value of the nickels and 10(30-3x) is the value of the dimes.........The total amount is given as 335 cents

OpenStudy (radar):

The equation will now look like this: 25x + 10x + 10(30-3x)=335 simplifying 25x +10 x +300 - 30x = 335 5x = 35 can you solve for the number of quarters now ?

OpenStudy (radar):

solving for x you would divide both sides by 5 getting x = 7 Or 7 quarters. 2x = 2 (7) = 14 nickels and 30 - 3(7)=30-21=9 dimes.

OpenStudy (anonymous):

OMG THANK YOU SO MUCH!!! I LOVE YOU YOUR A LIFESAVER!

OpenStudy (radar):

To verify 7 + 14 + 9 = 30 which is the number of coin now check the money. 1.75 + .70 + .90 = $3.35 They check out Your are welcome and good luck counting your money.

OpenStudy (radar):

Are their any questions? Did you follow with understanding?

OpenStudy (anonymous):

yes! love you man